Find the measures of the angles of an isosceles triangle if the measure of the vertex angle is 40 degrees less thean the sum of the measures of the base angles.
Answer by tutorcecilia(2152)

180=Angle #1+ Angle#2 + Angle#3 [Angle measures of a triangle]
Two of the angles of an isoceles triangle are equal.
Angle #1 = x
Angle #2=x
Angle #3=vertex angle =(x+x-40)
.
So,
180=x+x+(x+x-40)
180=4x-40
180+40=4x-40+40
220=4x
55=x
.
180=55+55+(55+55-40) [Plug all of the values back into the orginal formula]
180=180 [checks out]